Secret:

Write the equation 2x^2 - 4x + 6 = 1 in standard form and then choose the value of "c." 5 3 2.5

Shadow:

Hello @Secret Do you know what standard form is?

Secret:

Um, not really...

Shadow:

\[Ax + By = C\]

Shadow:

So for example, if I were to have: \[3x^2 - 2= 10x + 6\] I would add 2 to both sides, then subtract 10x from both sides, to get: \[3x^2 - 10x = 8\]

Secret:

Ohh

Shadow:

The number without a variable, also called the constant, should be alone on one side of the equation.

Secret:

Oh

Shadow:

Got it?

Secret:

Ya, kinda

Secret:

So i would have 2x^2 - 12x + 18 = 13???

Shadow:

How did you get 12, 18, and 13?

Secret:

Oof i messed up wait

Secret:

XD I think i messed it up with a different eqation

Secret:

I confused my self

Shadow:

Yeah. Otherwise you would have multiplied the 2nd and 3rd term by 3, then slapped a 3 behind the last term, lol

Secret:

Oof

Secret:

I'm so confused now...

Shadow:

It's simple. Have the terms with x on one side, and then the constant on the other side. \[ 2x^2 - 4x + 6 = 1\] \[2x^2 - 4x + 6 - 6 = 1 - 6\] \[2x^2 - 4x = -5\]

Shadow:

@Secret Does that make sense?

Secret:

Ohhhh, ya it does

Secret:

So then what do you do next???

Shadow:

choose the value of "c."